# A quantitative analysis of Galilei’s observations of Jupiter satellites from the Sidereus Nuncius

### v1

@@ 4 Other observations in the Sidereus Nuncius
In Fig. 17, left, we show the drawing of Galilei of the iconic Pleiades star open cluster in the constellation of the Bull. On the right a modern image. The two patterns has been superimposed in the bottom part where we have attempted to maximize the overlap by tuning the relative scaling and rotation. The cluster extends for about 1.5 degrees while the field of view of the cannocchiale was likely about one fifth of that. This can explain why the two patterns show several systematic distortions as it could not be visible all together at once. Nevertheless it is easy to establish an unambiguous correspondence with the stars in the modern picture in almost all cases. The names and magnitudes of some of these stars are shown in the figure. Galileo could see stars up to magnitude almost 9 (i.e. 8.94 for HD2336). It is interesting to ask if the positions of some of the stars might really have changed in 400 years such that some differences might be real. Considering that the proper motions can arrive to 60 mas/year this would translate into an angular displacement of 24 as or about half a prime corresponding to about the size of he Jupiter disk. The contribution of this effect is hence very marginal.
Figure 17: Pleiades observation in the Sidereus Nuncius and comparison with modern images.
@@
Galilei also reports observations of the Orion constellation. He writes that originally he wanted to accurately map the entire constellation but the task was overwhelming due to the large amount of visible stars. He then reports the region of the Orion belt where, as for the Pleiades, he marks the “old” (visible by naked eye) stars with large markers. A comparison of his map and the one from Stellarium is visible in Fig. 18.
Figure 18: The Orion belt region.
These maps resemble a bit old topographic maps of the World where basic features are correct but a lot of deformations are visible. We can state that with respect to the Pleiades the correspondence is less accurate. It should be noted that the field of view of the cannocchiale (about 1/3 of a degree) is tiny with respect to the extension of the map.
@@
It is not clear what the head of Orion nebula might be. The natural expectation is that it refers to M42 whose nebular appearance is also at reach with naked eye, but it does not sit where the head of the Orion constellation is located. We have compared the sketch of Galilei with the group of stars located where the head of Orion sits but it seems unlikely that the sketch represents this region of the sky. More likely it shows the region just north of M42 as in Fig. 19. In this case the association is tentative and this should be considered as a reasonable guess. The stars within M42 also does not seem to match in a convincing way so the previous interpretation seems the most likely one.
Figure 19: The “head of Orion” nebula.
The comparison for the Beehive cluster or M44 is shown in Fig. 20. Also in this case the match is not evident so we present two possible interpretations. Galilei states that the pair of large stars are the two “little donkeys” (Asellos, currently known as $\gamma$ and $\delta$ Cancri) that are also visible by naked eye. In this case the comparison is the one of Fig. 20 a). Still it seems strange that stars are not so clustered in a small region as in reality. It could be related by the difficulty of representing accurately with the symbols the differences in brightness. It could even be that what Galilei represented is something more similar to the case b) i.e. that the two “little donkeys” might have in reality been other stars closer to the cluster. Also in this case it is rather hard to establish a one-to-one correspondence unlikely to the case of the Pleiades.
a)Refer to caption
b)Refer to caption
Figure 20: Beehive cluster comparison. Possible interpretations.
We also tried to compare the beautiful sketch of the first quarter Moon with some modern telescopic images. In this case the accuracy seems a bit worse than in the case of Jupiter. Galilei manages to convey very well the effect of shades at the terminator for craters and mountain ridges but the overall fidelity of the representation in terms of dimensions and arrangement is not at the same level. The crater in the bottom half along the terminator has a significantly overestimated size. It might correspond to the the big crater Ptolemaeus as Tycho is smaller and more south. The big mountain arc could correspond to the Alps-Caucasus-Apennines mountains which are located between the Sea of Serenity and the Sea of Showers. In the sketch of Galilei one can recognize the Sea pattern given by Sea of Serenity, Tranquillity, Fecundity and Crises and Nectar. The mountain arc seems to surround the Sea of Serenity instead of being in between the Sea of Serenity and the Sea of Showers as in reality.
Figure 21: Comparison of the Moon sketches with modern pictures.
5 Conclusions
In this work we have analyse the observations of the satellites of Jupiter reported by Galileo Galilei in the Sidereus Nuncius [2] during the period from January 7 to March 1, 1610.
We compare them to the predictions obtained using a modern sky simulator [3], verifying them one by one. A sinusoidal fit of the data obtained from the 64 available sketches, allows measuring the relative major semi-axes of the satellites’ orbits and their periods with a statistical precision of 2-4% and 0.1-0.3% respectively.
The periods are basically unbiased while the orbits tend to be underestimated for Callisto by about 12%. The posterior fit error indicates that the positions of the satellites are determined with a resolution of 0.4-0.6 Jupiter diameters in the notation of Galilei corresponding to about 40-70 arc sec i.e. similar to the true angular diameter of Jupiter, in those days.
We show that with this data one can infer in a convincing way the third law of Kepler for the Jupiter system. The 1:2 and 1:4 Laplace orbital resonance between the periods of Io and Europa/Ganymede can be determined with % precision.
In order to obtain these results it is important to separate the four datasets. This operation, which is nowadays simple using a sky simulator, and is fully reported in this work, was an extremely difficult task for Galilei as the analysis will evidence. Nevertheless we show how the four periods might have been extracted using the modern Lomb-Scargle technique without having to separate the four data-sets already just using these early observations.
We also perform a critical evaluation of the accuracy of the observation of the Pleiades cluster, the beehive cluster, the Orion belt, the Orion’s head nebula and the Moon.
To conclude we have also used these data to infer the possible resolution and aberration of the “cannocchiale” of Galilei and use these parameters to imagine what could have been the view offered by Galilei to the head of the Serenissima republic (the “Doge”) when he was promoting his instrument from the Campanile of San Marco in Venice.
@@ 5 Conclusions
This analysis allows quantifying and appreciating the level of accuracy that Galilei managed to attain with the instrumentation he had at his disposal. Overall the accuracy of the measurements is quite impressing and adds to the value of a giant of early experimental physics.
@@ Appendix A Comparison of sketches and simulator
Figure 22: 1.
Figure 23: 2.
Figure 24: 3.
Figure 25: 4.
Figure 26: 5.
Figure 27: 6.
Figure 28: 7.
Figure 29: 8.
Figure 30: 9.
Figure 31: 10.
Figure 32: 11.
Figure 33: 12.
Figure 34: 13.
Figure 35: 14.
Figure 36: 15.
Figure 37: 16.
Figure 38: 17.
Figure 39: 18.
Figure 40: 19.
Figure 41: 20.
Figure 42: 21
Figure 43: 22
Figure 44: 23
Figure 45: 24
Figure 46: 25
Figure 47: 26
Figure 48: 27
Figure 49: 28
Figure 50: 29
Figure 51: 30
Figure 52: 31
Figure 53: 32
Figure 54: 33
Figure 55: 34
Figure 56: 35
Figure 57: 36
Figure 58: 37
Figure 59: 38
Figure 60: 39
Figure 61: 40
Figure 62: 41
Figure 63: 42
Figure 64: 43
Figure 65: 44
Figure 66: 45
Figure 67: 46
Figure 68: 47
Figure 69: 48
Figure 70: 49
Figure 71: 50
Figure 72: 51
Figure 73: 52
Figure 74: 53
Figure 75: 54
Figure 76: 55
Figure 77: 56
Figure 78: 57
Figure 79: 58
Figure 80: 59
Figure 81: 60
Figure 82: 61
Figure 83: 62
Figure 84: 63
Figure 85: 64
Appendix B Sunset correction
Figure 86: Sunset times in Padova as a function of the day since the beginning of the year. The red line shows a linear approximation that is used in the analysis to convert Galileo’s time data to the modern system as in Tab.  LABEL:tab:times.

### v2

@@ 6.1 The Pleiades
In Fig. 28, left, we show Sidereus Nuncius image of the iconic Pleiades star open cluster in the constellation of the Bull, the same where Jupiter was in those days. It was observed on Jan. 31. On the right a modern image. The two patterns has been superimposed in the bottom part where we have attempted to maximize the overlap by tuning the relative scaling and rotation. The cluster extends for about 1.5 while the field of view of the telescope was likely about one fifth of that. This can explain why the two patterns show several systematic distortions as it could not be visible all together at once. Nevertheless it is easy to establish an unambiguous correspondence with the stars in the modern picture in almost all cases. The names and magnitudes of some of these stars are shown in the figure.
Galileo could see stars up to magnitude almost 9 (i.e. 8.94 for HD2336). This is quite remarkable considering the small aperture of the telescope and it was helped by the total absence of light pollution in Padua, unlike nowadays.
It is interesting to wonder if the positions of some of the stars might really have changed in 400 years such that some differences might be real. Considering that the proper motions can arrive to 60 m′′/year this would translate into an angular displacement of 24 arcsec or about half a prime corresponding to about half the size of the Jupiter disk. The contribution of this effect is hence sub-leading (about half prime on a field of view of about 90 primes).
Figure 28: Pleiades: comparison with modern positions. The magnitudes are marked besides the stars. The faintest shines at 8.94.
6.2 The Orion belt and head
@@
Galilei also reports observations of the Orion constellation. He writes that originally he wanted to accurately map the entire constellation but the task was overwhelming due to the large amount of visible stars. He then reports the region of the Orion belt where, as for the Pleiades, he marks the “old” (visible by naked eye) stars with large markers. A comparison of his map and the one from Stellarium is in Fig. 29.
Figure 29: The Orion belt region. Red and yellow circles mark tentative associations between the sketch and the picture.
These maps resemble a bit old topographic maps of the World where basic features are correct but a lot of deformations are visible. We can state that with respect to the Pleiades the correspondence is less accurate. It should be noted that the field of view of the cannocchiale (about 1/3 of a degree) is tiny with respect to the extension of the map (represented as the white circle in the mid-top of the figure). It is interesting to notice that the rightmost big star does not form a right angle with the belt. Indeed if we inspect the original notes [32] of Feb. 7 1610, by which the Sidereus images are derived, it is clear that this is an error done in the printing phase as in the note the star position is more accurate (Fig. 30). Also the “sword” region is much more more realistic in the notes. It is interesting to notice the presence of two arcs connecting the belt and sword region that might indicate the fact that the sword is drawn closer than it should have been.
Figure 30: Comparison between the Orion belt in the Sidereus Nuncius (left) and on the original notes.
6.2.1 Absolute angular separations around Alnitak
In the same page of the note Galilei also reports angular separations between Alnitak (the leftmost star of the belt) and the surrounding stars. These data are rather clean as there are no uncertainties related to ephemerides and the precise knowledge of the observation time. Furthermore in this case there is no disk of Jupiter hence these data are not relative.
We have hence tried to check the accuracy of these angular measurements as shown in Fig. 31. In the left picture the original sketch is shown (Alnitak is the big star on the left). We have then taken the sketch and drawn the surrounding star following the written separations of 12, 30, 32, 45, 20 and 20 (red markers in the right plot) keeping the orientations of the original drawing. The blue markers represent modern values. A clear match emerges even though the azimuthal position correspondence is rather approximate. The distance of the closest star, (12, close to the maximal elongation of Callisto) is almost perfectly measured. The distances of the couple of stars at 20 are overestimated by few %. At larger distances (30-32) the overestimation becomes at the level of 20-30%. The separation of the farthest star at 45 (top-right) is also overestimated but less than before.
This analysis supports a general overestimation of the elongations even if these occur on a scale larger than the typical elongations of the satellites of Jupiter. The closest star is quite well matched, and the agreement is way better than the 30% discrepancy seen for the Jovian system. In this case, the overestimation at larger scale, likely comes from the need to work out of a single field of view. This might be the dominant effect.
It is worth noticing that Galileo reports in the same sheet a similar interesting sketch with distances of eight stars labelling it Canis. We tried to match it with some obvious objects: Sirius (Canis maior), Procyon (Canis minor) or the M41 open cluster (in Canis maior) or the near-by M47 open cluster but, unfortunately, we did not obtain a satisfactory match.
Figure 31: Alnitak and the surrounding stars with annotations on the angular separation (left). The right plot compares the modern angular distances with the ones on the left. The stars, shown by blue markers, are Alnitak (center), HD 37641 (mag. 7.53) the rightmost around $y=0$. Then, in clockwise order we find: HD37662 (mag. 8.5), HD37661 (mag. 8.8), V901-Ori (mag. 6.94), HD37903 (mag. 7.81), HD37805 (mag. 7.56) and HD37699 (mag. 7.56).
6.3 The Praesepe
@@
We have compared the sketch of Galilei with the group of stars located where the head of Orion sits. Despite being called a nebula this part of the constellation has nothing to do with “the sword” of Orion where the well-known Orion Nebula M42 lies. M42 was observed only later by Peiresc in 1610 and independently one year later by Cysat while Galilei later made a sketch of the stars of the trapezium asterism lying inside the M42 nebula ([33]). In Fig. 32 we have attempted a match with the simulator starfield. The bottom asterism (red circles) exhibits a quite convincing match while the rest of stars pattern has some resemblance but a massive distorsion effect must be assumed. Also in this case, the area is much larger than the field of view but more similar to the case of Pleiades than to the one of Orion. It is hence hard to understand why such a big distorsion effect is needed. For an independent evaluation it is interesting to mention an interesting work by Francesco Palla [33] done in 2009 where he also reaches similar conclusions using, in addition, real pictures taken with a replica of Galilei’s telescope and installed at the Arceteri observatory.
Figure 32: The head of Orion “nebula”.
The comparison for the Beehive cluster or M44 is shown in Fig. 33. In the following pictures the white circle highlights a field of view of 20. Galilei write that the pair of “large” stars are the two “little donkeys” (Asellos, $\gamma$ and $\delta$ Cancri), well visible by naked eye. What is striking is the absence of the little cluster of stars. The stars look more diffuse in the drawing. It is rather hard to establish a one-to-one correspondence unlikely to the case of the Pleiades. We have highlighted some tentative associations but there are several stars marked as bright that are hardly fitted in. The top image is the one that should correspond to the correct orientation where north is up. We have also considered, in the second image, flipping the drawing vertically (N-S). Other combinations, assuming other potential mistakes in the printing process, do not seem to give convincing matches. Also considering that the two brighter stars might not be the Asellos but two bright stars closer to the cluster does not bring much further.
Figure 33: Beehive cluster (M44) comparison. The top figure shows the drawing in its original form while the bottom one is flipped vertically.
6.4 Inkwashes of the Moon
A vast literature has discussed the interpretation and dating of the images of the moon [19, 34, 15, 35, 36, 37]. The engravings in the Sidereus Nuncius are based on seven inkwashes that are the most faithful representation of what Galilei saw. In the engravings a large crater was added to strengthen the message that the moon was an object similar to Earth. A discussion of the debates on this subject are given by Gingerich and Van Helden in 2023 in [19]. The accepted dates of the inkwashes are between 30 November 1609 and 17 December 1609 with the addition of an image on January 19 1610. It is remarkable that for this inkwash we can know the moment with resolution of minutes thanks to the vicinity of a star ($\theta$-Librae) that had just been reemerged after having been occulted by the Moon [38]. This happened at 6.36 AM CEST. Galileo, as we saw, had observed Jupiter also the previous evening (18/1) and he also did the following evening (19/1). Despite the urgency of the observations of Jupiter he had found the time to continue the observations of the Moon that he had started the month before.
Figures 38, 39, 40 (appendix) show a comparison of the seven inkwashes with the expectations of Stellarium. The hour and date were tuned to maximise the correspondence which, overall, is quite good. It is interesting to notice that in the last picture of Fig. 39, Galilei draws the Grimaldi crater a a grey small spot near the eastern limb of the waxing moon. Another interesting fact to observe is that the position of Mare Crisium is not matching so well, it seems much more extreme to the limb due to maximal libration in the simulation while in the inkwashes (especially n.3 and 4) it appears to be more central.
7 Tests with a modern replica of the telescope
[…]
Figure 34: Replica of the telescope.
Figure 35 is meant to convey the extremely narrow field of view by showing an example of the view provided for a terrestrial target in full daylight (left and central pictures). The central picture shows that there are no easily identifiable distortions near the edge of the field. The right picture shows the Jupiter system pictured on 21 Jan. 2025 and compared to Stellarium. The quality of the picture is rather poor as it was simply taken with a smartphone in a lucky shot. Still the presence of stray light is quite evident as well as the glare of the Jupiter disk. It is not difficult to imagine the difficulty of spotting satellites near the disk. But what is the real lesson coming from a direct experience with this object is the need for a really good support. On one hand even touching the object with the eye creates a risk for losing the target. In addition the target drifts very fast [^6] It is easy to calculate that the target moves from the center to the edge in about 45 s with a f.o.v. of 20. in the field of view and thus the possibility to apply fine corrections is needed. Pointing Jupiter might require some minutes. Callisto can easily fall out of the field of view. Initially it was tested on a support for cameras but later the equatorial mount of a modern telescope was used. It must also be stressed that, as we have seen, Galilei had to operate it with altitude angles up to about 70. The idea of using the micrometer techniques inputs additional challenges.
These challenges are not mentioned by Galilei in the Sidereus but he mentions them in a letter of 7 January 1610 to Antonio De Medici (or Enea Piccolomini) when he says ([4], pag. 62): Che lo strumento si tenga fermo, et perciò è bene, per fuggire la titubatione della mano che dal moto delle arterie et della respiratione stessa essa procede, fermare il cannocchiale in qualche luogo stabile. i.e. “That the instrument be held steady, and therefore it is advisable, in order to avoid the unsteadiness of the hand — which arises from the motion of the arteries and from breathing itself — to fix the telescope upon some stable support” [4]. Bonifacio Vannozzi mentions the activity of Sergio Venturi, that in Rome was trying to build telescopes in that period between 1609 and 1610 and says: ([4], pag. 59) But in the sea, where these instruments could be very beneficial, they can never be used well, for the continuous motion of that setting. Another indication of the frustration in the use of these instruments is found in a letter of Giovanni Bartoli to Belisario Vinta ([4], pag. 33) As for me, having seen several of them — and in particular one that was sold for three zecchini to the master of the post in Prague — I must confess I am not fully satisfied with it, because, being longer than an arm’s length, one has to struggle for quite a while to catch with the eye the object one wishes to see, and once found, one must hold the instrument so steady that the slightest movement makes it vanish again.
The difficulty of the observation must have been particularly frustrating when Galilei went to Bologna to convince his rival Magini and a quite numerous group of “well-educated people” about his observations. Magini writes to Kepler[4], pag. 91: On the 24th and 25th of April he stayed at my house with his telescope, wishing to show the new satellites of Jupiter. He did not succeed. In fact, more than twenty most learned men were present, yet not one of them was able to clearly see the new planets.
Figure 35: View of the galileian telescope view during daylight on a terrestrial target. Notice the narrowness of the field of view. Observation of the Galileian satellites with the modern reproduction of the Galileian telescope. Taken 21 Jan 2025 at 19.14 CEST.
Figure  36 shows a collage of pictures of the Moon taken with the replica on Feb. 4, 2025. A significant chromatic aberration can be noticed in the southern part were the craters are visible. Near the border some vignetting is visible (also in the central picture of Fig. 35).The angular dimension of the Moon was 32.5 from which we can infer a field of view of about 15. Overall this picture testifies that the performance of the original was probably not much worse than the replica. The inkwashes presented in Sec. 6.4 are definitely a good job considering the drift of the field of view, its limited extension, the need to realign, the vibrations and so on.
Figure 36: Collage of pictures of the Moon taken with the reproduction of the Galileian telescope. Pictures of the moon taken on February 4 2025 at about 20 CEST from Padua. In the rightmost one Mare Crisium is visible. In the topo one the craters Exodus and Aristoteles (from bottom to top). In the second from the top Mare Serenitatis with Appennines and Caucasus at the terminator. In the southern region the two biggest visible craters are Maurolycus and Gemma Frisius.
We plan to use this replica to involve students in a study attempting to reproduce the Sidereus dataset and compare their performances with those achieved by Galilei.
8 Summary and conclusions
In this work we have extend significantly the understanding of the Sidereus Nuncius dataset by applying an in depth analysis of the full achievable data:
  • we have presented a detailed comparison of two datasets (Sec. 22.1): the one coming from an analysis of the drawings in the published Sidereus and the angular data given by Galilei in the text. We show how digitized data suffer less from discretization and clarify how the angular measurements must be referred to the limb in order to make sense. To reconcile positive and negative elongations it is necessary to assume a Jupiter disk of 67-79′′, as obtained from a comparison of the two sets, using a $\chi^{2}$ minimization. The angular data do not evidence a convincing variation of the apparent angle of the Jovian system that is expected at the 15% level. This hints to the fact that these are not really absolute angular measurements but still relative to Jupiter, where its angular size is assumed equal to 1. Despite discussing a way to do an absolute calibration of the measurement Galilei seems to get the angles wrong, in excess with respect to the real ones. The normalization to the angular size of Jupiter, that is overestimated with respect to the real dimension, could also explain the overestimation of the elongations with respect to the ephemerides (about 70% lower than the data). We performed a check of some angular measurements taken in the original notes on stars around Alnitak. In this case a separation of 12 is estimated quite correctly, without a 30% discrepancy.
  • We have presented sinusoidal fits of the elongations for both datasets as a function of the modern time of the 64 (dataset-1) or 64-5+9 (dataset-2) observations. The fitted orbital parameters are already extremely close to modern expectations. The errors are at sub-percent level for periods and at few % level for the orbit’s radiuses. By normalizing to the orbit of Ganymede we find that the elongations are accurate at better than 5% for Io and Europa while Callisto is underestimated by about 5-10%, depending on the considered dataset.
  • The residuals of the fit are used to estimate the resolution of the measurements. The errors range from 0.5-0.7 Jupiter diameters in the sketches or (0.86-1.44).
  • For the first time we have shown how a frequency analysis of sparse data (Lomb-Scargle periodograms) allows getting a fairly solid evaluation of the orbital periods of Callisto and Ganymede without having to disentangle the individual satellites;
  • We have extend the study proposed by Bettini to this dataset showing how the data provide a stringent test of the third Kepler law for the Jovian system and a clear indication of the 1:2:4 resonance of the periods of the innermost satellites (for Io this is possible only by using the simulator to isolate the data of each satellite).
  • We have evaluate the correctness of all the configurations and discussed the only known anomaly of Feb. 12. Each configuration drawn by Galilei was presented beside to the prediction from the software allowing to evaluate in a direct and intuitive way the reasons for missing observations.
  • We have reevaluated the resolving power of pairs of satellites and the inefficiency introduced by the proximity to Jupiter’s disk by analyzing the distribution of the positions of the missed satellites or missed pairs at “truth-level” using the simulator. We have shown rigorously that the onset of the efficiency drop starts already at a distance of about three diameters with respect to the center of Jupiter and falls rapidly to zero. The closest observations are typically pairs of satellites that were not resolved and hence resulted as a single satellite with double luminosity thus compensating the difficulty of detection close to the planet. We have created a 2D map of inefficient observations and of the positions of unresolved satellite pairs whose separation is typically around 15-17′′.
  • In Sec. 6 we have found an incredible level of accuracy for the observation of the Pleiades while for the Beehive cluster, the Orion head and belt, the association is more problematic, very likely due to the smallness of the field of view and/or time limits.
  • The seven inkwashes of the Moon as predicted for the widely accepted dates show a relatively good match with the simulator in terms of phases and recognizable details; here we have not embarked on considerations on the pattern of Maria to infer information about the libration of the Moon.
[…]
@@ 8 Summary and conclusions
This analysis allows quantifying and appreciating the level of accuracy that Galilei managed to attain with the instrumentation he had at his disposal. Even more after trying a replica of the telescope, the achieved accuracy and the diligence and completeness of these (early!) measurements is impressing and it can only add to the value of Galilei as a giant of (experimental) physics.
Acknowledgements: I would like to express my gratitude to Alessandro Bettini, Carlo Broggini, Giulio Peruzzi, Enrico Maria Corsini and Mauro Mezzetto of the Department of Physics and Astronomy “Galileo Galilei” and INFN Padova for their encouraging comments on an early version of this work and guidance to the relevant literature.
@@ Appendix A Comparison of sketches and simulator
Table 10: Side-by-side images with progressive numbering
No. Sidereus Nuncius Simulation
1 [Uncaptioned image] [Uncaptioned image]
2* [Uncaptioned image] [Uncaptioned image]
3* [Uncaptioned image] [Uncaptioned image]
4* [Uncaptioned image] [Uncaptioned image]
5 [Uncaptioned image] [Uncaptioned image]
6* [Uncaptioned image] [Uncaptioned image]
7 [Uncaptioned image] [Uncaptioned image]
8 [Uncaptioned image] [Uncaptioned image]
9 [Uncaptioned image] [Uncaptioned image]
10 [Uncaptioned image] [Uncaptioned image]
11 [Uncaptioned image] [Uncaptioned image]
12 [Uncaptioned image] [Uncaptioned image]
13 [Uncaptioned image] [Uncaptioned image]
14 [Uncaptioned image] [Uncaptioned image]
15 [Uncaptioned image] [Uncaptioned image]
16 [Uncaptioned image] [Uncaptioned image]
17 [Uncaptioned image] [Uncaptioned image]
18 [Uncaptioned image] [Uncaptioned image]
19 [Uncaptioned image] [Uncaptioned image]
20 [Uncaptioned image] [Uncaptioned image]
21 [Uncaptioned image] [Uncaptioned image]
22 [Uncaptioned image] [Uncaptioned image]
23* [Uncaptioned image] [Uncaptioned image]
24 [Uncaptioned image] [Uncaptioned image]
25 [Uncaptioned image] [Uncaptioned image]
26 [Uncaptioned image] [Uncaptioned image]
27 [Uncaptioned image] [Uncaptioned image]
28 [Uncaptioned image] [Uncaptioned image]
29 [Uncaptioned image] [Uncaptioned image]
30 [Uncaptioned image] [Uncaptioned image]
31 [Uncaptioned image] [Uncaptioned image]
32 [Uncaptioned image] [Uncaptioned image]
33* [Uncaptioned image] [Uncaptioned image]
34 [Uncaptioned image] [Uncaptioned image]
35 [Uncaptioned image] [Uncaptioned image]
36 [Uncaptioned image] [Uncaptioned image]
37 [Uncaptioned image] [Uncaptioned image]
38* [Uncaptioned image] [Uncaptioned image]
39* [Uncaptioned image] [Uncaptioned image]
40 [Uncaptioned image] [Uncaptioned image]
41 [Uncaptioned image] [Uncaptioned image]
42 [Uncaptioned image] [Uncaptioned image]
43 [Uncaptioned image] [Uncaptioned image]
44 [Uncaptioned image] [Uncaptioned image]
45 [Uncaptioned image] [Uncaptioned image]
46 [Uncaptioned image] [Uncaptioned image]
47 [Uncaptioned image] [Uncaptioned image]
48* [Uncaptioned image] [Uncaptioned image]
49 [Uncaptioned image] [Uncaptioned image]
50 [Uncaptioned image] [Uncaptioned image]
51 [Uncaptioned image] [Uncaptioned image]
52 [Uncaptioned image] [Uncaptioned image]
53* [Uncaptioned image] [Uncaptioned image]
54 [Uncaptioned image] [Uncaptioned image]
55 [Uncaptioned image] [Uncaptioned image]
56 [Uncaptioned image] [Uncaptioned image]
57 [Uncaptioned image] [Uncaptioned image]
58 [Uncaptioned image] [Uncaptioned image]
59 [Uncaptioned image] [Uncaptioned image]
60 [Uncaptioned image] [Uncaptioned image]
61 [Uncaptioned image] [Uncaptioned image]
62 [Uncaptioned image] [Uncaptioned image]
63 [Uncaptioned image] [Uncaptioned image]
64* [Uncaptioned image] [Uncaptioned image]
Appendix B Measurements without sketches
Figure 37: Comparison between the simulator and the observations with only angular data but no associated sketches. The top sketch is built from the angular separations by Galilei. Bottom image the prediction.
Appendix C Moon inkwashes
Figure 38: Comparison of the Moon inkwashes with the prediction from the Stellarium simulator. The date is shown in the upper left side of the picture. The white line marks the terminator in the inkwash and it is superimposed to image from the simulation in the left to guide the eye in the comparison.
Figure 39: Comparison of the Moon inkwashes with the prediction from the Stellarium simulator. The date is shown in the upper left side of the picture. The white line marks the terminator in the inkwash and it is superimposed to image from the simulation in the left to guide the eye in the comparison.
Figure 40: Comparison of the Moon inkwashes with the prediction from the Stellarium simulator. The date is shown in the upper left side of the picture. The white line marks the terminator in the inkwash and it is superimposed to image from the simulation in the left to guide the eye in the comparison.
Appendix D Conversion from Italic hours
January 1610 February–March 1610
id mo. day italic (h) sunset modern id mo. day italic (h) sunset modern
1 1 7 1:00 16.74 17:44 32 2 1 2:00 17.33 19:19
2 1 8 1:00 16.77 17:46 33 2 2 1:00 17.36 18:21
3 1 10 1:00 16.81 17:48 34 2 2 7:00 17.36 24:21
4 1 11 1:00 16.84 17:50 35 2 3 7:00 17.38 24:22
5 1 12 1:00 16.86 17:51 36 2 4 2:00 17.40 19:24
6 1 13 1:00 16.89 17:53 37 2 4 7:00 17.40 24:24
7 1 15 3:00 16.93 19:55 38 2 6 1:00 17.45 18:26
8 1 15 7:00 16.93 23:55 39 2 7 1:00 17.47 18:28
9 1 16 1:00 16.96 17:57 40 2 8 1:00 17.50 18:29
10 1 17 0:30 16.98 17:28 41 2 9 0:30 17.52 18:01
11 1 17 5:00 16.98 21:58 42 2 10 1:30 17.54 19:02
12 1 18 0:20 17.00 17:20 43 2 11 1:00 17.57 18:34
13 1 19 2:00 17.03 19:01 44 2 11 3:00 17.57 20:34
14 1 19 5:00 17.03 22:01 45 2 11 5:30 17.57 23:04
15 1 20 1:15 17.05 18:17 46 2 12 0:40 17.59 18:15
16 1 20 6:00 17.05 23:02 47 2 13 0:30 17.61 18:06
17 1 20 7:00 17.05 24:02 48 2 15 1:00 17.66 18:39
18 1 21 0:30 17.07 17:34 49 2 15 5:00 17.66 22:39
19 1 22 2:00 17.10 19:05 50 2 15 6:00 17.66 23:39
20 1 22 6:00 17.10 23:05 51 2 16 6:00 17.68 23:41
21 1 23 0:40 17.12 17:47 52 2 17 1:00 17.71 18:42
22 1 23 5:00 17.12 22:07 53 2 18 1:00 17.73 18:43
23 1 24 1:00 17.14 18:08 54 2 18 6:00 17.73 23:43
24 1 24 6:00 17.14 23:08 55 2 19 0:40 17.76 18:25
25 1 25 1:40 17.17 18:50 56 2 21 1:30 17.80 19:18
26 1 26 0:40 17.19 17:51 57 2 25 1:30 17.90 19:23
27 1 26 5:00 17.19 22:11 58 2 26 0:30 17.92 18:25
28 1 27 1:00 17.21 18:12 59 2 26 5:00 17.92 22:55
29 1 30 1:00 17.29 18:17 60 2 26 5:00 17.92 22:55
30 1 31 2:00 17.31 19:18 61 2 27 1:04 17.94 19:00
31 1 31 4:00 17.31 21:18 62 2 28 1:00 17.97 18:58
63 2 28 5:00 17.97 22:58
64 3 1 0:40 17.99 18:39
Table 11: Conversion from Italic to modern hours (January–March 1610).
Figure 41: Comparison of the absolute predictions from the ephemerides and the fit for dataset-2 with an ad-hoc scaling factor on the amplitude of about 0.7.